An animal shelter has five times as many cats as dogs. Find the total number of cats if the total number of cats and dogs is 48.
A. 5 cats
B. 8 cats
C. 13 cats
D. 40 cats

Answer by jim_thompson5910(35256)

c+d=48

c=5d


c+d=48

5d+d=48

6d = 48

d = 48/6

d = 8

c = 5d

c = 5(8)

c = 40

So there are 40 cats